OpenStudy (anonymous):

Rationalize the denominator: 4 / 7- square root of 6 ..........@SolomonZelman ??

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{4}{7-\sqrt{6}} }\) For this we have a `conjugate`. For a denominator ` a - √b` it is `a + √b`. Knowing the rule `(a-b)(a+b)=a²-b²` you can see that there is no middle term, since it cancels out, right? So for your denominator will do the same.

OpenStudy (solomonzelman):

Multiply top and bottom times ` 7 + √6 ` . let me know if you need more help.

OpenStudy (anonymous):

Thank you!!!!!! Makes a lot of sense, i kept getting confused

OpenStudy (solomonzelman):

Can we clarify what you get? Tell me your final answer please :)